# Efficient Estimation of the Central Mean Subspace via Smoothed Gradient Outer Products

### v1

@@ 3.5 Instantiation for Polynomial Links
$\tau:=\min_{\boldsymbol{\eta}\in\mathbb{R}^{k}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0}_{k},\boldsymbol{I}_{k})}\left[\left(\boldsymbol{\eta}^{{\scriptscriptstyle{\mathsf{T}}}}\nabla f(\boldsymbol{Z})\right)^{2}\right]\,.$
@@
The fact that $\tau\neq 0$ is guaranteed by the minimality of $\mathcal{U}$, since otherwise, for some $\boldsymbol{\eta}\in\mathbb{R}^{k}$, $\boldsymbol{U}\boldsymbol{\eta}\in\mathcal{U}$ would be an irrelevant direction for the regression function $g$.
@@
Suppose that 6 holds. Let $\widehat{\boldsymbol{U}}$ be the output from Algorithm 1 with inputs $20<h<\sqrt{5}/$ 2, $0<\sigma_{\theta}<h/\sqrt{20}$ and $m=n/(2d)$. Then, there exists $C_{2}>0$, independent of $n$, $d$, $\delta$, $h$ and $\sigma_{\theta}$, such that with probability at least $1-\delta$,
@@
where $A_{h,d,r},B_{\sigma_{\theta}^{2}/h^{2},d,r}$ are given as:
@@
The dependence on $d$ in the bound of Corollary 3.19 is hard to improve since our estimator falls in the class of so-called Correlational Statistical Query (CSQ) learners, i.e., learners that rely solely on statistics of form $\sum_{i}Y_{i}\cdot\Phi(\boldsymbol{X}_{i})$ for some function $\Phi$ (see [2, 18, 13] for an exact definition). Namely, such CSQ learners have been shown in the recent work of [12] to require a sample complexity $n=\Omega(d^{r/2})$ to achieve nontrivial error in subspace estimation. In comparison, our method requires $n=O(d^{2r})$, which corresponds to a constant factor mismatch in the exponent of $d$. It remains unclear how such a gap may be closed for CSQ learners as the result of [12] also leaves this open. We note however that, outside of CSQ procedures, much more benign sample size requirements are possible, for instance $n=O(d)$ was shown for a recent method of [5].
@@ 4 Analysis Overview
$=:\sum_{\boldsymbol{\alpha}\in\mathbb{N}^{k}}\lambda_{\boldsymbol{\alpha},i}S_{\boldsymbol{\alpha}}(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{\theta})\,,$ (8)
@@
$\forall\boldsymbol{\alpha}\in\mathbb{N}^{k},\quad\sum_{i\in[k]}\eta_{i}\lambda_{\boldsymbol{\alpha},i}=0\,.$ (9)
@@
$=\sum_{\boldsymbol{\alpha}\in\mathbb{N}^{k}}\lambda_{\boldsymbol{\alpha},i}\cdot H_{\boldsymbol{\alpha}}(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{x})\,.$ (10)
@@
^$\lVert\widehat{}\boldsymbol{M}-\widetilde{\boldsymbol{M}}\rVert_{\operatorname{op}}\leq C\left(\tilde{C}_{h,\sigma_{\theta},1}\cdot d(\mu_{f}+\sigma_{Y})^{2}\cdot\frac{\sqrt{m}}{n}+\tilde{C}_{h,\sigma_{\theta},2}\cdot d^{1/2}(\mu_{f}+\sigma_{Y})\cdot\frac{1}{\sqrt{n}}\right)\frac{\log({4d}/{{\delta}})}{\sqrt{{\delta}}}\,.$ (11)
@@ 5 Simulations
In this section, we demonstrate how to choose the hyperparameters $h$ (the “radius” of smoothing) and $m$ (the number of $\boldsymbol{\theta}_{j}$ sampled in ASGOP) properly through a simple example with polynomial link function and standard Gaussian design. The related theoretical discussions can be found in Remark 3.5, 3.10, 3.15 and 3.20.
@@ Appendix A SAVE Is Not Exhaustive
$Y=f(\boldsymbol{u}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X})\,,$ (12)
where $\boldsymbol{X}=(X_{1},X_{2})^{{\scriptscriptstyle{\mathsf{T}}}}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{2})$, and $\boldsymbol{u}=(1,0)^{{\scriptscriptstyle{\mathsf{T}}}}$. Next, we show that there is a class of link functions with which $\boldsymbol{V}$ is a zero matrix. Let $\xi(z_{1},z_{2}):=\operatorname*{\mathbb{E}}_{Z\sim\mathcal{N}(0,1)}[Z^{2}|Z\in[z_{1},z_{2}]]$ for $0\leq z_{1}\leq 1\leq z_{2}$. Note that $\xi$ is continuous, and monotonically increasing in each of $z_{1}$ and $z_{2}$, $\xi(1,1)=\operatorname*{\mathbb{E}}[Z^{2}|Z=1]=1$ and $\xi(0,\infty)=\operatorname*{\mathbb{E}}[Z^{2}|Z\geq 0]=1$. Therefore, there exists a continuous decreasing function $\nu:[1,\infty)\to(0,1]$, such that $\nu(1)=1,\lim_{z\to\infty}\nu(z)=0$ and
@@
$f(z):=\left\{\begin{array}[]{lll}0,&&z<0;\\ f_{0}(z),&&0\leq z\leq 1;\\ f_{0}(\nu(z)),&&1<z<\infty.\end{array}\right.\,$ (16)
@@ Appendix B Connection to Local Linear Regression
$\hat{\boldsymbol{\beta}}^{\operatorname{LLE}}_{h}(\boldsymbol{\theta})=\boldsymbol{I}_{d}^{\backslash 1}(\boldsymbol{X}_{\boldsymbol{\theta}}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{W}^{\boldsymbol{\theta}}_{h}\boldsymbol{X}_{\boldsymbol{\theta}})^{-1}\boldsymbol{X}_{\boldsymbol{\theta}}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{W}^{\boldsymbol{\theta}}_{h}\boldsymbol{Y}$ (18)
@@ Appendix C Estimation Bias Under $P_{\boldsymbol{X}}$ with a Bounded Support
$=-{h^{-2}}\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[\mathbbm{1}(\lVert\boldsymbol{Z}\rVert>K)\cdot g(\boldsymbol{Z})\cdot(\boldsymbol{Z}-\boldsymbol{\theta}_{j})\right]$ (19)
Intuitively, $\lVert B_{h,j}\rVert$ should shrink to zero as $h\rightarrow 0$ when $K\gg\lVert\boldsymbol{\theta}_{j}\rVert$, which holds with high probability if we choose the variance $\sigma_{\theta}^{2}$ of $\boldsymbol{\theta}_{j}$’s is small compared to $K$. On the other hand, it is clear that if $\sup_{j\in[m]}\lVert B_{h,j}\rVert=o(n^{-1/2})$, then the bias will not affect the rate in the upper-bound for subspace estimation error. The following proposition shows how small $h$ needs to be to guarantee such small biases.
@@
$\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[g(\boldsymbol{Z})^{3}\right]\leq C\cdot\mu_{\rho}\cdot\mu_{f}\,.$ (20)
@@
$\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[\lVert\boldsymbol{Z}-\boldsymbol{\theta}_{j}\rVert^{3}\right]\leq C\cdot h^{3}\,.$ (21)
@@
$\operatorname*{\mathbb{E}}_{\widetilde{\boldsymbol{Z}}\sim\mathcal{N}(h^{-2}\boldsymbol{\theta}_{j},\boldsymbol{I}_{d})}\left[\mathbbm{1}(\lVert\widetilde{\boldsymbol{Z}}\rVert>Kh^{-1})\right]$ $\leq\operatorname*{\mathbb{P}}\left(\lVert\mathcal{N}(\boldsymbol{0}_{d},\boldsymbol{I}_{d})\rVert>\frac{Kh^{-1}}{2}\right)=\operatorname*{\mathbb{P}}\left(\chi^{2}_{d}>\frac{K^{2}h^{-2}}{4}\right).$ (22)
@@
$\operatorname*{\mathbb{P}}\left(\chi^{2}_{d}>\frac{K^{2}h^{-2}}{4}\right)\leq n^{-2}\,.$ (23)
@@
The condition $\lVert\boldsymbol{\theta}_{j}\rVert\leq h$ in Proposition C.1 holds uniformly for $j\in[m]$, with probability at least $1-\delta$ with the choice $\sigma_{\theta}\lesssim h\cdot\log(m/\delta)$. This is compatible with our choice of $\sigma_{\theta}$ in the main theorem.
See pages - of supplementary.pdf

### v2

@@ 3.5 Instantiation for Polynomial Links
$\tau:=\min_{\boldsymbol{\eta}\in\mathbb{S}^{k-1}}\ \operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0}_{k},\boldsymbol{I}_{k})}\left[\left(\boldsymbol{\eta}^{{\scriptscriptstyle{\mathsf{T}}}}\nabla f(\boldsymbol{Z})\right)^{2}\right]\,,$
where $\mathbb{S}^{k-1}\doteq\{\boldsymbol{x}\in\mathbb{R}^{k}:\lVert\boldsymbol{x}\rVert=1\}$ is the sphere of the $k$-dimensional unit ball.
@@
The fact that $\tau\neq 0$ is guaranteed by the minimality of $\mathcal{U}$, since otherwise, for some $\boldsymbol{\eta}\in\mathbb{S}^{k-1}$, $\boldsymbol{U}\boldsymbol{\eta}\in\mathcal{U}$ would be an irrelevant direction for the regression function $g$.
@@
Suppose that 6 holds.
Let $\widehat{\boldsymbol{U}}$ be the output from Algorithm 1 with inputs $20<h<\sqrt{5}/$ 2, $0<\sigma_{\theta}<h/\sqrt{20}$ and $m=n/(2d)$. Then, there exists $C_{2}>0$, independent of $n$, $d$, $\delta$, $h$ and $\sigma_{\theta}$, such that with probability at least $1-\delta$,
@@
where $A_{\sigma_{\theta}^{2}/h^{2},d,r},B_{h,d,r}$ are given as:
@@
The dependence on $d$ in the bound of Corollary 3.20 is hard to improve since our estimator falls in the class of so-called Correlational Statistical Query (CSQ) learners, i.e., learners that rely solely on statistics of form $\sum_{i}Y_{i}\cdot\Phi(\boldsymbol{X}_{i})$ for some function $\Phi$ (see [2, 20, 15] for an exact definition). Such CSQ learners have been shown in the recent work of [13] to require a sample complexity $n=\Omega(d^{r/2})$ to achieve nontrivial error in subspace estimation with bounded adversarial noises. As a heuristic comparison, our method requires $n=O(d^{2r})$, which corresponds to a constant factor mismatch in the exponent of $d$. A recent work [14] shows that such a gap can be closed for single-index models with the knowledge of the link function. However, for general multi-index models, it remains unclear how such a gap may be closed for CSQ learners as the result of [13] also leaves this open.
We note however that, outside of CSQ procedures, much more benign sample size requirements are possible, for instance $n=O(d)$ was shown for a recent method of [6].
@@ 4 Analysis Overview
(8) $=:\sum_{\boldsymbol{\alpha}\in\mathbb{N}^{k}}\lambda_{\boldsymbol{\alpha},i}S_{\boldsymbol{\alpha}}(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{\theta})\,,$
@@
(9) $\forall\boldsymbol{\alpha}\in\mathbb{N}^{k},\quad\sum_{i\in[k]}\eta_{i}\lambda_{\boldsymbol{\alpha},i}=0\,.$
@@
(10) $=\sum_{\boldsymbol{\alpha}\in\mathbb{N}^{k}}\lambda_{\boldsymbol{\alpha},i}\cdot H_{\boldsymbol{\alpha}}(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{x})\,.$
@@
(11) ^$\lVert\widehat{}\boldsymbol{M}-\widetilde{\boldsymbol{M}}\rVert_{\operatorname{op}}\leq C\left(\tilde{C}_{h,\sigma_{\theta},1}\cdot d(\mu_{f}+\sigma_{Y})^{2}\cdot\frac{\sqrt{m}}{n}+\tilde{C}_{h,\sigma_{\theta},2}\cdot d^{1/2}(\mu_{f}+\sigma_{Y})\cdot\frac{1}{\sqrt{n}}\right)\frac{\log({4d}/{{\delta}})}{\sqrt{{\delta}}}\,.$
@@ 5 Simulations
In this section, we demonstrate how to choose the hyperparameters $h$ (the “radius” of smoothing) and $m$ (the number of $\boldsymbol{\theta}_{j}$ sampled in ASGOP) properly through a simple example with polynomial link function and standard Gaussian design. The related theoretical discussions can be found in Remark 3.5, 3.11, 3.16 and 3.21.
@@ Appendix A SAVE Is Not Exhaustive
(12) $Y=f(\boldsymbol{u}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X})\,,$
where $\boldsymbol{X}=(X_{1},X_{2})^{{\scriptscriptstyle{\mathsf{T}}}}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}_{2})$, and $\boldsymbol{u}=(1,0)^{{\scriptscriptstyle{\mathsf{T}}}}$. Next, we show that there is a class of link functions with which $\boldsymbol{V}$ is a zero matrix. Let $\xi(z_{1},z_{2}):=\operatorname*{\mathbb{E}}_{Z\sim\mathcal{N}(0,1)}[Z^{2}|Z\in[z_{1},z_{2}]]$ for $0\leq z_{1}\leq 1\leq z_{2}$. Note that $\xi$ is continuous, and monotonically increasing in each of $z_{1}$ and $z_{2}$, $\xi(1,1)=\operatorname*{\mathbb{E}}[Z^{2}|Z=1]=1$ and $\xi(0,\infty)=\operatorname*{\mathbb{E}}[Z^{2}|Z\geq 0]=1$.
Therefore, there exists a continuous decreasing function
$\nu:[1,\infty)\to(0,1]$, such that $\nu(1)=1,\lim_{z\to\infty}\nu(z)=0$ and
@@
(16) $f(z):=\left\{\begin{array}[]{lll}0,&&z<0;\\ f_{0}(z),&&0\leq z\leq 1;\\ f_{0}(\nu(z)),&&1<z<\infty.\end{array}\right.\,$
@@ Appendix B Connection to Local Linear Regression
(18) $\hat{\boldsymbol{\beta}}^{\operatorname{LLE}}_{h}(\boldsymbol{\theta})=\boldsymbol{I}_{d}^{\backslash 1}(\boldsymbol{X}_{\boldsymbol{\theta}}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{W}^{\boldsymbol{\theta}}_{h}\boldsymbol{X}_{\boldsymbol{\theta}})^{-1}\boldsymbol{X}_{\boldsymbol{\theta}}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{W}^{\boldsymbol{\theta}}_{h}\boldsymbol{Y}$
@@ Appendix C Estimation Bias Under $P_{\boldsymbol{X}}$ with a Bounded Support
(19) $=-{h^{-2}}\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[\mathbbm{1}(\lVert\boldsymbol{Z}\rVert>K)\cdot g(\boldsymbol{Z})\cdot(\boldsymbol{Z}-\boldsymbol{\theta}_{j})\right]$
Intuitively, $\lVert B_{h,j}\rVert$ should shrink to zero as $h\rightarrow 0$ when $K\gg\lVert\boldsymbol{\theta}_{j}\rVert$, which holds with high probability if we choose the variance $\sigma_{\theta}^{2}$ of $\boldsymbol{\theta}_{j}$’s is small compared to $K$.
On the other hand, it is clear that if $\sup_{j\in[m]}\lVert B_{h,j}\rVert=o(n^{-1/2})$, then the bias will not affect the rate in the upper-bound for subspace estimation error. The following proposition shows how small $h$ needs to be to guarantee such small biases.
@@
(20) $\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[g(\boldsymbol{Z})^{3}\right]\leq C\cdot\mu_{\rho}\cdot\mu_{f}\,.$
@@
(21) $\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta}_{j},h^{2}\boldsymbol{I}_{d})}\left[\lVert\boldsymbol{Z}-\boldsymbol{\theta}_{j}\rVert^{3}\right]\leq C\cdot h^{3}\,.$
@@
(22) $\operatorname*{\mathbb{E}}_{\widetilde{\boldsymbol{Z}}\sim\mathcal{N}(h^{-2}\boldsymbol{\theta}_{j},\boldsymbol{I}_{d})}\left[\mathbbm{1}(\lVert\widetilde{\boldsymbol{Z}}\rVert>Kh^{-1})\right]$ $\leq\operatorname*{\mathbb{P}}\left(\lVert\mathcal{N}(\boldsymbol{0}_{d},\boldsymbol{I}_{d})\rVert>\frac{Kh^{-1}}{2}\right)=\operatorname*{\mathbb{P}}\left(\chi^{2}_{d}>\frac{K^{2}h^{-2}}{4}\right).$
@@
(23) $\operatorname*{\mathbb{P}}\left(\chi^{2}_{d}>\frac{K^{2}h^{-2}}{4}\right)\leq n^{-2}\,.$
@@
The condition $\lVert\boldsymbol{\theta}_{j}\rVert\leq h$ in Proposition C.1 holds uniformly for $j\in[m]$, with probability at least $1-\delta$ with the choice $\sigma_{\theta}\lesssim h\cdot\log(m/\delta)$. This is consistent with our choice of $\sigma_{\theta}$ in the main theorem.
Appendix D The Case with Unknown Density Ratio
In this section, we consider the estimation of the smoothed gradient when the density ratio is unknown. Without loss of generality, we assume that the smoothing radius $h=1$ throughout this section and omit all $h$ in the subscript and simply write $\boldsymbol{\beta}_{j}:=\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})$. Recall that for $j\in[m]$, $\ell=1,2$, we have the following smoothed gradient estimators in Section 3.3:
$\hat{\boldsymbol{\beta}}_{j,l}=\frac{2m}{n}\sum_{(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j}),\quad$
One natural solution to resolve the issue of unknown $\rho$ is to replace it by some estimates $\hat{\rho}_{N}$ based on another sample ~$\widetilde{}\boldsymbol{X}_{1},\ldots,\widetilde{}\boldsymbol{X}_{N}$, and we get the following plug-in estimator:
$\tilde{\boldsymbol{\beta}}_{j,\ell}=\frac{2m}{n}\sum_{(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j}).$
The following proposition quantifies the effect of “plugging in” on the $L_{2}$ error of estimating the smoothed gradients.
Proposition D.1.
Let $j\in[m]$, $\ell=1,2$, and $\boldsymbol{x},\boldsymbol{\theta}_{j}\in\mathbb{R}^{d}$. Suppose that $\hat{\rho}_{N}(\boldsymbol{x};\boldsymbol{\theta}_{j})$ is an estimate of the density ratio $\rho(\boldsymbol{x};\boldsymbol{\theta}_{j})$ from a sample ~$\widetilde{}\boldsymbol{X}_{1},\ldots,\widetilde{}\boldsymbol{X}_{N}$ with:
$\operatorname*{\mathbb{E}}\left[\left|\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})\right|^{12}\right]<\infty\,,\operatorname*{\mathbb{E}}\left[\lVert\boldsymbol{X}-\boldsymbol{\theta}_{j}\rVert^{12}\right]<\infty\,,$
$a_{N,j}:=\operatorname*{\mathbb{E}}\left[\left(\frac{\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})}{\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)^{4}\right]\overset{N\to\infty}{\longrightarrow}1,$
$b_{N,j}:=\operatorname*{\mathbb{E}}\left[\left({\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})-\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)^{2}\right]\overset{N\to\infty}{\longrightarrow}0,$
where the expectation is taken over the randomness of $\boldsymbol{X},\widetilde{\boldsymbol{X}}_{1},\ldots,\widetilde{\boldsymbol{X}}_{N}$. Then, we have for some $C>0$ independent of $n,m$ and $N$ such that:
$\operatorname*{\mathbb{E}}\left[\lVert\tilde{\boldsymbol{\beta}}_{j,\ell}-\boldsymbol{\beta}_{j}\rVert^{2}\right]\leq C\cdot\left(\frac{m}{n}\cdot a_{N,j}^{n/4m}+b_{N,j}\right).$
Remark D.2.
The performance of the plug-in estimator ~$\tilde{}\boldsymbol{\beta}_{j,\ell}$ depends on both the relative error $a_{N,j}$ and the absolute error $b_{N,j}$ of the density ratio estimator $\hat{\rho}_{N}(\boldsymbol{x};\boldsymbol{\theta}_{j})$. The same squared $L_{2}$ error rate of $O(m/n)$ as the oracle estimator ^$\hat{}\boldsymbol{\beta}_{j,\ell}$ can be obtained when $a_{N,j}^{n/4m}$ is bounded from above by a constant and $b_{N,j}$ is of smaller order than $O(m/n)$.
Proof D.3.
We first apply the standard $L_{2}$ error decomposition:
$\operatorname*{\mathbb{E}}[\lVert\tilde{\boldsymbol{\beta}}_{j,\ell}-{\boldsymbol{\beta}}_{j}\rVert^{2}]$ $=\operatorname*{\mathbb{E}}[\lVert\tilde{\boldsymbol{\beta}}_{j,\ell}-\operatorname*{\mathbb{E}}\tilde{\boldsymbol{\beta}}_{j,\ell}\rVert^{2}]+\lVert\operatorname*{\mathbb{E}}{\boldsymbol{\beta}}_{j,\ell}-{\boldsymbol{\beta}}_{j}\rVert^{2}$
(24) $\leq\operatorname*{\mathbb{E}}[\lVert\tilde{\boldsymbol{\beta}}_{j,\ell}\rVert^{2}]+\lVert\operatorname*{\mathbb{E}}{\boldsymbol{\beta}}_{j,\ell}-{\boldsymbol{\beta}}_{j}\rVert^{2}.$
In what follows, we provides upper bound the two terms on the right-hand side.
$\operatorname*{\mathbb{E}}[\lVert\tilde{\boldsymbol{\beta}}_{j,\ell}\rVert^{2}]$ $=\operatorname*{\mathbb{E}}\left[\lVert\hat{\boldsymbol{\beta}}_{j,\ell}\rVert^{2}\cdot\prod_{\boldsymbol{X}:(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\left(\frac{\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})}{{\rho}(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)^{2}\right]$
$\leq\left(\operatorname*{\mathbb{E}}[\lVert\hat{\boldsymbol{\beta}}_{j,\ell}\rVert^{4}]\right)^{1/2}\cdot\left(\operatorname*{\mathbb{E}}\left[\left(\frac{\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})}{{\rho}(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)^{4}\right]\right)^{n/4m}$
$=\left(\operatorname*{\mathbb{E}}\left[\left\lVert\frac{2m}{n}\sum_{(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j})\right\rVert^{4}\right]\right)^{1/2}\cdot a_{N}^{n/4m}$
$\leq\frac{2m}{n}\cdot\left(\operatorname*{\mathbb{E}}\left[\left\lVert\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j})\right\rVert^{4}\right]\right)^{1/2}\cdot a_{N}^{n/4m}$
$\leq\frac{2m}{n}\cdot\left(\operatorname*{\mathbb{E}}\left[\left|\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})\right|^{12}\right]\operatorname*{\mathbb{E}}\left[\lVert\boldsymbol{X}-\boldsymbol{\theta}_{j}\rVert^{12}\right]\operatorname*{\mathbb{E}}\left[Y^{12}\right]\right)^{1/6}\cdot a_{N}^{n/4m}$
$\lesssim\frac{2m}{n}\cdot a_{N,j}^{n/4m}\,.$
$\lVert\operatorname*{\mathbb{E}}\tilde{}\boldsymbol{\beta}_{j,\ell}-{\boldsymbol{\beta}}_{j}\rVert$ $=\lVert\operatorname*{\mathbb{E}}[\left({\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})-\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j})\cdot Y]\rVert$
$\leq\left(\operatorname*{\mathbb{E}}[\left({\hat{\rho}_{N}(\boldsymbol{X};\boldsymbol{\theta}_{j})-\rho(\boldsymbol{X};\boldsymbol{\theta}_{j})}\right)^{2}]\right)^{1/2}\cdot\left(\operatorname*{\mathbb{E}}[Y^{4}]\right)^{1/4}\cdot\left(\operatorname*{\mathbb{E}}[\lVert\boldsymbol{X}-\boldsymbol{\theta}_{j}\rVert^{4}]\right)^{1/4}$
$\lesssim b_{N,j}^{1/2}\,.$
The proof is completed by substituting these bounds in D.3.
Appendix E Improving Dependence on $\delta$
Recall that in the upper-bound of our main theorem (Theorem 3.14), the dependence on failure probability $\delta$ is of order $\tilde{O}(\delta^{-1/2})$ which may appear large in light of usual high-probability results, e.g., sub-gaussian concentrations. This is mainly due to the heavy-tailedness of the outer product estimates ^$\widehat{}\boldsymbol{M}_{j}$’s. We note that such dependence on $\delta$ is further improvable by the classical idea of median-of-mean estimators in robust statistics [26]. Here, we present Algorithm 2, a median-of-mean variant of our estimator, adapted from Algorithm 2 of [19]. The main difference from Algorithm Algorithm 1 lies in the way of aggregating the ASGOP estimates $\{\widehat{\boldsymbol{M}}_{j}\}_{j\in[m]}$. Instead of taking a simple average, the multivariate median ^$\widehat{}\boldsymbol{M}_{\star}$ (c.f. Line 7) is taken as the final estimate of the ASGOP $\widetilde{\boldsymbol{M}}$. It can be readily shown that with proper choice of $m\sim\log(1/\delta)$, one can establish an exponential concentration of ^$\widehat{}\boldsymbol{M}_{\star}$ to $\widetilde{\boldsymbol{M}}$. Here, we outline the key steps:
  • By LABEL:lem:Z-error,
    $P(\lVert\widehat{}\boldsymbol{M}_{j}-\widetilde{\boldsymbol{M}}\rVert\geq C_{d}\cdot\sqrt{m/n})\leq 2/3\,,$
    for some $C_{d}>0$, which may depends on dimension $d$ and various parameters defined as in Section 3. Hence, Assumption 1 and 2 in [19] is satisfied with $\varepsilon=C_{d}\cdot\sqrt{m/n}$.
  • By Proposition 9 of [19], we have:
    $P\left(\lVert\widehat{}\boldsymbol{M}_{\star}-\widetilde{\boldsymbol{M}}\rVert\geq 3C_{d}\cdot\sqrt{m/n}\right)\leq e^{-m/8}\,.$
  • Choose $m\approx 8\log(1/\delta)$ to establish the desired results.
Algorithm 2 Estimation of the CMS $\mathcal{U}$ (Median-of-Means Variant)
2:Randomly sample $\{\boldsymbol{\theta}_{j}\}_{j=1}^{m}$ from $\mathcal{N}(\mathbf{0}_{d},\sigma_{\theta}^{2}\mathbf{I}_{d})$
5:$\forall j\in[m]$, let $\hat{\boldsymbol{\beta}}_{j,1}\leftarrow\hat{\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j},\mathcal{D}_{j,1})$, and $\hat{\boldsymbol{\beta}}_{j,2}\leftarrow\hat{\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j},\mathcal{D}_{j,2})$triangleright See (3.4) in the main paper
6:$\forall j\in[m]$, ^$\widehat{}\boldsymbol{M}_{j}\leftarrow\frac{1}{2}(\hat{\boldsymbol{\beta}}_{j,1}\hat{\boldsymbol{\beta}}^{{\scriptscriptstyle{\mathsf{T}}}}_{j,2}+\hat{\boldsymbol{\beta}}_{j,2}\hat{\boldsymbol{\beta}}^{{\scriptscriptstyle{\mathsf{T}}}}_{j,1})$; and {^${\cal M}\leftarrow\{\widehat{}\boldsymbol{M}_{j}:j\in[m]\}$
7:$\forall j\in[m]$, $r_{j}:=\min\{r\geq 0:|B_{r}(\widehat{}\boldsymbol{M}_{j})\cap{\cal M}|>m/2\}$; $j_{\star}:=\arg\min_{j\in[m]}r_{j}$; and ^$\widehat{}\boldsymbol{M}_{\star}\leftarrow\widehat{}\boldsymbol{M}_{j_{\star}}$
8:Return: $\widehat{\mathcal{U}}_{\star}\in\mathbb{R}^{d\times k}$, the top-$k$ eigenvectors of ^$\widehat{}\boldsymbol{M}_{\star}$
See pages - of ESGOP_supplementary.pdf